OpenStudy (anonymous):

What is the least common multiple,LCM, of the denominators 3/x+2+2/x=1?

OpenStudy (anonymous):

\[\frac{ 3 }{ x+2 }+\frac{ 2 }{ x }= 1\]

OpenStudy (ivettef365):

between (x+2) and x would be x(x+2)

OpenStudy (anonymous):

Soo is the answer just x(x+2)?

OpenStudy (zzr0ck3r):

yep